8. Project A has a required return on 9.2 percent and cash flows of −$87,000, $32,600, $35,900, and $43,400 for Years 0 to 3, respectively. Project B has a required return of 12.7 percent and cash flows of −$85,000, $14,700, $21,200, and $89,800 for Years 0 to 3, respectively. Which project(s) should you accept based on net present value if the projects are mutually exclusive?

Project B. Please type out all work as I am on mobile please no excel or grid style work
Answer #1

Project A:

CF0 = -87,000

Cf1 = 32,600

CF2 = 35,900

CF3 = 43,400

Required rate = 9.2%

NPV of Project A = -87,000 + 32,600/ (1 + 9.2%)1 + 35,900/ (1 + 9.2%)2 + 43,400/ (1 + 9.2%)3

NPV of Project A = -87,000 + 32,600/ 1.092 + 35,900/ 1.192 + 43,400/ 1.302

NPV of Project A = -87,000 + 29,853.48 + 30,105.73 + 33,328.96

NPV of Project A = 6,288.17

Project B:

CF0 = -85,000

Cf1 = 14,700

CF2 = 21,200

CF3 = 89,800

Required rate = 12.7%

NPV of Project B = -85,000 + 14,700/ (1 + 12.7%)1 + 21,200/ (1 + 12.7%)2 + 89,800/ (1 + 12.7%)3

NPV of Project B = -85,000 + 14,700/ 1.13 + 21,200/ 1.27 + 89,800/ 1.43

NPV of Project B = -85,000 + 13,043.48 + 16,691.22 + 62,734.23

NPV of Project B = 7,468.93

As NPV of project B is higher we will choose project B